OpenStudy (anonymous):

Find the cube roots of 27(cos 330° + i sin 330°

OpenStudy (anonymous):

@agent0smith please help :(

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

have they covered De Moivre's theorem yet?

OpenStudy (anonymous):

Yes i just dont get it

OpenStudy (ranga):

\[\Large e^{ix} = \cos x + isin x\]

OpenStudy (anonymous):

r^3(cos(3Θ + 360nº)+i sin(3Θ + 360nº)) z^3= 27(cos 330º + i sin 330º) z1 = 3(cos 110º + i sin 110º) z2 = 3(cos 230º + i sin 230º) z3 = 3(cos 350º + i sin 350º) i have that

OpenStudy (anonymous):

So is what i have the answer or do i need to do more?

OpenStudy (ranga):

\[\Large (27(\cos 330 + isin 330))^{1/3} = (27e^{i330})^{1/3} = 3(\cos 110 + i \sin 110)\]

OpenStudy (ranga):

No that is it.

OpenStudy (anonymous):

so should i earase all that i have

OpenStudy (ranga):

No, those are the steps. We start with the Euler's formula: e^(ix) = cos(x) + isin(x) Then to find cube root of 27(cos 330° + i sin 330°) we raise it to the power of 1/3: {27(cos 330° + i sin 330°)}^(1/3) = {27e^(330i)}^(1/3) = 3e^(110i) = 3(cos 110º + i sin 110º)

OpenStudy (anonymous):

Thank you~